at the same temperature which of the following gases contains molecules with the slowest average velocity?
A) NO2
B) CO
C) CO2
D) NO
E) the average velocity is the same
Answer #1

rms(root mean square) velocity of gases is defined by v_{rms}=\sqrt{\frac{3RT}{M}}

Most probable velocity v_{mp}=\sqrt{\frac{2RT}{M}}

Average velocity v_{av}=\sqrt{\frac{8RT}{\pi M}}

Where, M= molar mass of gas, T= temperature, R = universal gas constant

So at same temperature as the molar mass increases average velocity decreases.

M.W_{NO_2}=46\; g/mol

M.W_{CO}=28\; g/mol

M.W_{CO_2}=44\; g/mol

M.W_{NO}=30\; g/mol

So trend of molar mass is M.W_{CO} <M.W_{NO}<M.W_{CO_2}<M.W_{NO_2}

So trend of average speed will be v_{av}(CO)>v_{av}(NO)>v_{av}(CO_2)>v_{av}(NO_2)

Answer:

A) NO2 has slowest average velocity.